# Bridging Gaps in Natural Language Processing for Yorùbá: A Systematic Review of a Decade of Progress and Prospects

### v1

@@ 4.2.3 RQ3: What language resources are available for Yorùbá language?
YFACC: YFACC [35] connected speech with visual representations in the dataset comprising spoken captions in Yorùbá language for $6,000$ Flickr [^15] https://www.flickr.com images. The audio captions also utilised single-speaker recording to ensure consistency in the audio data. Moreover, it includes cross-lingual pairing, whereby images are tagged with English visual labels and paired with Yorùbá speech, permitting cross-lingual applications. The empirical analysis is also catered for by developing a baseline cross-lingual model. Ultimately, this dataset addressed the dearth of visually-grounded speech datasets in low-resource languages, essentially focusing on Yorùbá language. Furthermore, it provides a new benchmark for visually-grounded speech models in low-resource language settings.
Yorùbá Speech Corpus: An open-source speech dataset exclusively for Yorùbá developed in [36]. This corpus comprises over $4$ hours of $48kHz$ recordings from $36$ male and female volunteers. It also includes transcriptions with disfluency annotations and full diacritization, essential for pronunciation and lexical disambiguation. Moreover, the dataset was tested in a statistical parametric speech synthesis (SPSS) for evaluation purposes and compared with related corpora in the same domain. The corpus will support TTS, ASR, and speech-to-speech translation, contributing to West African corpus linguistics.
@@
YORÙLECT: YORÙLECT [34] introduces a high-quality parallel text and speech corpus specifically in Yorùbá language, across three domains—machine translation, automatic speech recognition, and speech-to-text translation—and four Yorùbá dialects— Standard Yorùbá, Ifẹ̀, Ìjẹ̀bú, and Ìlàjẹ. The Ifẹ̀ dialect is spoken primarily among the people of Ọ̀ṣun state, Ìjẹ̀bú among people of Ọ̀gùn state, and Ìlàjẹ, among the people of Oǹdó state—all states located in the South West geopolitical zone of Nigeria. Text data from various sources were obtained and localised into the three Yorùbá dialects for the corpus development. Similarly, high-quality speech data in standard Yorùbá, Ifẹ̀, and Ìlàjẹ were recorded. The text portion contains about $1506$ parallel sentences for each dialect, totalling $6024$, and the speech part contains $\approx 3$ hours of audio for the trio of standard Yorùbá, Ifẹ̀, and Ìlàjẹ. Furthermore, to aid empirical analysis, experiments involving the domain of text MT, ASR and speech-to Yorùbá were carried out, and the dataset and models were published under an open licence, making it relevant for Yorùbá NLP tasks.
Table LABEL:tab7 is used to summarize the available language resources relevant for NLP development involving Yorùbá language.
@@
AFRIQA Annotated corpus Cross-lingual open-retrieval question answering Multilingual [18]
AfriSenti Annotated corpus Sentiment analysis Multilingual [4]
@@
CIRAL Annotated corpus Cross-lingual information retrieval Multilingual [111]
@@
MasakhaNER Annotated corpus Named entity recognition Multilingual [12]
@@
MIRACL Annotated corpus Monolingual information retrieval Multilingual [109], [108]
@@
NaijaSenti Annotated corpus Sentiment analysis Multilingual [6]
@@ 4.2.4 RQ4: What are the major challenges in developing NLP solutions for Yorùbá?
Natural language solution development for Yorùbá faces multiple challenges limiting its growth as a low-resource language. Data synthesis in this study has helped identify five primary challenges, including linguistic, technical, resource, cultural and societal factors, and evaluation and benchmarking challenges.
@@
Linguistic challenges: Yorùbá language exhibits several linguistic properties that make NLP development in it an exigent task. These linguistic features of the language often complicate its NLP development, and they mainly account for its tone dependency, morphology, diacritic demand, and code-switching demerits, which are summarised briefly below:
@@
  • Tonal complexity: Yorùbá language implicit pitch contours, denoting tonal components, in communicating emphasis and other para-linguistic expressions. Thus, they usually determine the semantic property of the message conveyed in the language. For instance, Ògún (a deity associated with iron) is different from either of ogun (war) and ogún (twenty). This distinction in linguistic meaning complicates various NLP tasks even though the words belong to the same phonetic sequence. For instance, one of the main challenges for speech recognition in Yorùbá is determining the tone associated with a syllable [91]. Similarly, utilising context-dependent phone units to capture long-term spectrum dependencies of tone in Yorùbá is typically less successful and oftentimes requires a different means of acoustic modelling [93]. Furthermore, in bilingual English-Yorùbá MT tasks, tone-changing verbs [66] also present challenges as they frequently alter the semantic properties of the sentence they are used in by shifting from a low-tone to a mid-tone. Consequently, accurate representation and processing of tones are essential for NLP tasks like diacritic restoration, TTS synthesis, and MT.
@@
  • Diacritic dependency: Similar to the tonal feature of Yorùbá language, it also requires appropriate assignment of accent property to characters in a segment. This task is usually referred to as diacritic restoration, and it’s essential to fully decipher the linguistic meaning of words. Diacritic marks are usually placed above, below, or between characters to indicate pronunciation and may change the meaning of the composed words [53]. Like most languages involving diacritic, Yorùbá language users often omit them in electronic texts, increasing lexical ambiguity and also posing challenges to NLP systems as a result of mislaying the accompanying syntactic, grammatical, or semantic information, partially or totally [51]. This omission is sometimes due to the unavailability of supporting applications and devices [31] or lack of knowledge by most users [90], resulting in a drawback towards improving NLP representation in the language.
@@
  • Morphological Complexity: Carrying out analysis of word formation and structure in Yorùbá language tends to be oftentimes complex due to its plenitude in terms of rules involving noun and verb inflection patterns [58]. Handling agglutinative morphology and word compounding has also been stated to require advanced morphological analysis tools [59]. The deficiency in this regard frequently creates ambiguity in unsophisticated NLP systems due to issues like affixes of words and the required rules to correctly programme the system for effective analysis. Efforts towards addressing this have incorporated automatic morphological induction [57] to ensure compatibility, producing a more accurate representation in the NLP domain.
@@
  • Code-Switching: This is common for most Nigerian languages, and Yorùbá is not spared of the growing ‘civilization’. Despite initiatives towards sustaining native languages, it is evident that the younger generation finds it difficult to carry on lengthy conversations or even to form lengthy sentences without interspersing the conversation with terms frequently borrowed from the English language [30]. This phenomenon introduces additional complexity in NLP tasks like language identification, sentiment analysis, and machine translation, as they reduce data availability purely in the language. This is evident in the dataset obtained for NaijaSenti [6], which aimed to conduct monolingual sentiment analysis for four languages through a curated corpus from X, as a certain proportion of the corpus had to cater for the phenomenon.
@@
Resource challenges: The challenges in obtaining substantial or quality resources for training NLP models involving Yorùbá language is the ultimate pitfall for their NLP development, just like in many other low-resource languages [76, 72]. The existence of a comprehensive lexicon for most high-resource languages makes their pre-processing task less taxing, unlike a low-resource language like Yorùbá. The unavailability often leads to applying manual data cleaning, hence the task of token validation for corpus development [125].
@@
In addition, limited corpora pose significant challenges in various NLP tasks. For instance, a comparatively short corpus of speech recordings and minimal language-specific development can automatically build acoustic models for an elementary speech synthesiser in a new language; however, tonal languages like Yorùbá require additional resources [102] which is usually not readily available. Also, word-level and character-level models are especially common for most diacritic restoration tasks [53]. However, they require significant training data to prevent sparsity in the models. Similarly, out-of-vocabulary word translation is a significant issue for low-resource languages that lack parallel training data [70]. This phenomenon is considered a stumbling block to NLP development in these different domains, albeit recent efforts have been directed towards implementing other viable methods requiring lower resources for a handful of tasks supporting such.
@@
Generally, limited available corpora for developing NLP tools have been a consistent challenge in NLP development involving Yorùbá and under low-resource languages. Studies have also shown the presence of quality issues for existing corpora and models [76], which limits the reliability of findings from using such corpus. Consequently, efforts towards creating readily available corpora for specific NLP tasks are sacrosanct since better results mostly require high-quality and large data [51].
@@
Other language tools, including the syntactic parser and POS tagger, have also been reported to mostly fail in accounting for Yorùbá language linguistic nuances [15], thereby constituting a challenge requiring the development of language-specific tools for these specific tasks. Similarly, an essential NLP task such as effective spell-checking development in Yorùbá is still at the early development [56] due to non-consideration of diacritic necessity in earlier studies [55], resulting in their limitation. A limited number of studies [17, 54, 56] involving computational NLP have been seen to include Yorùbá language in these domains. Furthermore, the majority of Yorùbá texts are also seen to be typed using the plain ASCII Character Set, without diacritics [31], perhaps due to limited tools supporting their full implementation. This situation also induces lower quality translation between Yorùbá language and most European languages due to lack of adequate elision resolution tool [30]. Thus, there is a growing need for flexible, multilingual pre-processing solutions.
Cultural and Societal Factors: An African language, such as Yorùbá, is not only a mirror into the mind of its users but also a mirror into their culture and history [38]. This emphasises the richness of the culture and language. However, younger generations of Yorùbá speakers increasingly favour English for formal and digital communication, while the parents also decide not to teach their infants many times [90]. This shift reduces the volume of contemporary Yorùbá texts and contributes to the decline of the language in digital contexts. Moreover, despite the evident progress in MT tasks for low-resource languages, NMT models still lag in accurately carrying out automatic translation involving cultures [72] for Yorùbá language. Consequently, cultural and societal challenges of Yorùbá in NLP development require considerable attention, as they could potentially endanger the language in the face of westernization, globalization, and inter-ethnic communication [38].
@@ 5 Discussion
This systematic review was intended to capture the growth and current stage of NLP participation for Yorùbá language over a decade. Through the information synthesised from the primary studies, it is obvious that significant progress has been made in notable areas of NLP involving Yorùbá language. MT has particularly received more attention through the inclusion of the language primarily in bilingual MT research [64, 68, 74, 69, 28, 67, 50] involving translation between two languages. Related studies investigating the inherent prerequisite towards enhancing translation qualities in MT model qualities, the development of MT tools, such as vowel elision resolution [30], and the application of rough set theory [29] in translation have been explored. Moreover, the need to create better quality data [76] has also led towards improved research in this domain through the development of additional parallel corpora and leveraging multilingual language pre-training to improve the model training capability in the presence of large high-quality datasets.
Similar to MT research, significant efforts have been made in TTS synthesis, with considerable efforts towards corpus development, as evidenced by publicly available datasets, most of which are speech corpora. Even though it is the primary focus in most research investigating the task, some of the studies have included ASR and speech-to-text [101, 97, 34] in their experiments contributing to the entire research. Furthermore, sentiment analysis and information retrieval tasks involving Yorùbá have also benefited significantly from multilingual corpus development through the need to develop language resources for low-resource African languages. Most have been through the works in [4] and [111].
@@
Furthermore, recent studies have showcased the importance of multilingual and cross-lingual techniques, as they help to promote language availability for many African languages simultaneously. This is evident from named entity recognition [12, 14], sentiment analysis [6, 4, 5], and information retrieval tasks [18, 111, 109]. Also, models pre-trained on multilingual datasets have been seen to be beneficial for a low-resource language like Yorùbá. Ultimately, it is also evident that collaborative, open-source and community initiatives have greatly improved research and availability of resources in the bid to overcome resource scarcity whilst fostering knowledge-sharing among these resource-scarce African languages.
@@
However, in the context of constraints stunting the rapid growth of NLP in Yorùbá language, studies have highlighted mainly linguistic complexity such as diacritic dependency [31] and tonal variation [66]. Other challenges involving cultural and societal factors were also highlighted, including the primary hindrance, which has always been a limitation in available corpora or the quality of such available corpora.
5.1 Limitation of study

### v2

@@ 4.2.3 RQ3: What language resources are available for Yorùbá language?
YFACC: YFACC [96] connected speech with visual representations in the dataset comprising spoken captions in Yorùbá language for $6,000$ Flickr [^15] https://www.flickr.com images. The audio captions also utilised single-speaker recording to ensure consistency in the audio data. Moreover, it includes cross-lingual pairing, whereby images are tagged with English visual labels and paired with Yorùbá speech, permitting cross-lingual applications. The empirical analysis is also catered for by developing a baseline cross-lingual model. Ultimately, this dataset addressed the dearth of visually-grounded speech datasets and a new benchmark in low-resource languages, specifically the Yorùbá language.
Yorùbá Speech Corpus: This contains an open-source speech dataset exclusively for Yorùbá [65]. This corpus comprises over $4$ hours of recordings from $36$ male and female volunteers. It also includes transcriptions with disfluency annotations and full diacritization, essential for pronunciation and lexical disambiguation. Moreover, the dataset was tested in a statistical parametric speech synthesis (SPSS) for evaluation purposes and compared with related corpora in the same domain. The corpus supports TTS, ASR, and speech-to-speech translation, contributing to West African corpus linguistics.
@@
YORÙLECT: YORÙLECT [22] introduces a high-quality parallel text and speech corpus specifically in Yorùbá language, across three domains—machine translation, automatic speech recognition, and speech-to-text translation—and four Yorùbá dialects—Standard Yorùbá, Ifẹ̀, Ìjẹ̀bú, and Ìlàjẹ. The Ifẹ̀ dialect is spoken primarily among the people of Ọ̀ṣun state, Ìjẹ̀bú among people of Ọ̀gùn state, and Ìlàjẹ, among the people of Oǹdó state; all states located in the South West geopolitical zone of Nigeria. Text data from various sources were obtained and localised into the three Yorùbá dialects for the corpus development. Similarly, high-quality speech data in standard Yorùbá, Ifẹ̀, and Ìlàjẹ were recorded. The text portion contains about $1506$ parallel sentences for each dialect, totalling $6024$, and the speech part contains $\approx 3$ hours of audio for the considered dialects. Furthermore, to aid empirical analysis, experiments involving the domains of MT, ASR, and speech-to-text were done. The dataset and models were published under an open licence, making it relevant for Yorùbá NLP tasks.
Table LABEL:tab7 is used to summarize the available language resources relevant for NLP development involving the Yorùbá language.
@@
AFRIQA Annotated text corpus Cross-lingual open-retrieval question answering Multilingual [89]
AfriSenti Annotated text corpus Sentiment analysis Multilingual [85]
@@
CIRAL Annotated text corpus Cross-lingual information retrieval Multilingual [19]
@@
MasakhaNER Annotated text corpus Named entity recognition Multilingual [10]
@@
MIRACL Annotated text corpus Monolingual information retrieval Multilingual [128], [94]
@@
NaijaSenti Annotated text corpus Sentiment analysis Multilingual [86]
@@
Synthesis and Implications of Available Resources
While the listed resources reflect commendable progress in Yorùbá NLP, there exist significant limitations that shape the scope of current research. For instance, although multiple corpora exist, most are either small-scale, domain-specific, or lack detailed linguistic annotation, which hinders their use for advanced syntactic or semantic modeling tasks. The absence of standard benchmarks further complicates reproducibility and model comparison.
Moreover, lexical and morphological resources such as POS taggers primarily support rule-based or traditional machine learning methods. In contrast, resources suited for training modern transformer-based models—such as large-scale annotated datasets or pretrained embeddings specifically for Yorùbá—are sparse or underutilized. In addition, limited speech resources restrict progress in speech-to-text or conversational systems.
These gaps suggest that while foundational resources exist, their fragmented and often narrow nature limits scalability and generalization. Going forward, resource creation should further prioritize open licensing, broader domain coverage, and alignment with task-specific benchmarks to better support the evolving landscape of low-resource NLP.
@@ 4.2.4 RQ4: What are the major challenges in developing NLP solutions for Yorùbá?
The development of natural language solutions for Yorùbá faces multiple challenges, which hinder the language’s representation and accessibility in digital and computational contexts, especially as a low-resource language. Data synthesis in this study has helped identify five primary challenges, including linguistic, technical, resource, cultural, and societal factors, and evaluation and benchmarking challenges.
@@
Linguistic challenges: The Yorùbá language exhibits several linguistic properties that make NLP development in it an exigent task. These include its tone-dependence, complex morphology, diacritics dependency, and code-switching language usage. Each of these challenges is briefly described below:
@@
  • Tonal complexity: Yorùbá language’s implicit pitch contours, denoting tonal components, is used in communicating emphasis and other para-linguistic expressions. Thus, they usually determine the semantic property of the message conveyed in the language. For instance, Ògún (a deity associated with iron) is different from either of ogun (war) and ogún (twenty). This distinction in linguistic meaning complicates various NLP tasks even though the words belong to the same phonetic sequence. For instance, one of the main challenges for speech recognition in Yorùbá is determining the tone associated with a syllable [49]. Similarly, utilising context-dependent phone units to capture long-term spectrum dependencies of tone in Yorùbá is typically less successful and oftentimes requires a different means of acoustic modelling [115]. Furthermore, in bilingual English-Yorùbá MT tasks, tone-changing verbs [67] also present challenges as they frequently alter the semantic properties of the sentence they are used in by shifting from a low-tone to a mid-tone. Consequently, accurate representation and processing of tones are essential for NLP tasks like diacritic restoration, TTS synthesis, and MT.
@@
  • Diacritic dependency: Similar to the tonal feature of Yorùbá language, it also requires appropriate assignment of accent property to characters in a segment. This task is usually referred to as diacritic restoration, and it is essential to fully decipher the linguistic meaning of words. Diacritic marks are usually placed above, below, or between characters to indicate pronunciation and may change the meaning of the composed words [34]. Like most languages involving diacritics, Yorùbá language users often omit them in electronic texts, increasing lexical ambiguity and also posing challenges to NLP systems as a result of mislaying the accompanying syntactic, grammatical, or semantic information, partially or totally [44]. This omission is sometimes due to the unavailability of supporting applications and devices [101] or lack of knowledge by most users [1], resulting in a drawback towards improving NLP representation in the language.
@@
  • Morphological Complexity: Carrying out analysis of word formation and structure in Yorùbá language tends to be oftentimes complex due to its plenitude in terms of rules involving noun and verb inflection patterns [21]. Handling agglutinative morphology and word compounding has also been stated to require advanced morphological analysis tools [7]. The deficiency in this regard frequently creates ambiguity in unsophisticated NLP systems due to issues like affixes of words and the required rules to correctly program the system for effective analysis. Efforts towards addressing this have incorporated automatic morphological induction [8] to ensure compatibility, producing a more accurate representation in the NLP domain.
@@
  • Code-Switching: This is common for most Nigerian languages, and Yorùbá is not spared of the growing ‘civilization’. Despite initiatives towards sustaining native languages, it is evident that the younger generation finds it difficult to carry on lengthy conversations or even to form lengthy sentences without interspersing the conversation with terms frequently borrowed from the English language [16]. This phenomenon introduces additional complexity in NLP tasks like language identification, sentiment analysis, and machine translation, as it reduces data availability purely in the language. This is evident in the dataset obtained for NaijaSenti [86], which aimed to conduct monolingual sentiment analysis for four languages through a curated corpus from X, as a certain proportion of the corpus had to cater to the phenomenon.
@@
Resource challenges: The challenges in obtaining substantial or quality resources for training NLP models involving Yorùbá language are the ultimate pitfall for their NLP development, just like in many other low-resource languages [93, 13]. The existence of a comprehensive lexicon for most high-resource languages makes their pre-processing task less taxing, unlike a low-resource language like Yorùbá. The unavailability often leads to applying manual data cleaning, hence the task of token validation for corpus development [17].
@@
In addition, limited corpora pose significant challenges in various NLP tasks. For instance, a comparatively short corpus of speech recordings and minimal language-specific development can automatically build acoustic models for an elementary speech synthesiser in a new language; however, tonal languages like Yorùbá require additional resources [123], which is usually not readily available. Also, word-level and character-level models are especially common for most diacritic restoration tasks [34]; however, they require significant training data to prevent sparsity in the models. Similarly, out-of-vocabulary word translation is a significant issue for low-resource languages that lack parallel training data [80]. This phenomenon is considered a stumbling block to NLP development in these different domains, albeit recent efforts have been directed towards implementing other viable methods requiring lower resources for a handful of tasks supporting such.
@@
Generally, limited available corpora for developing NLP tools have been a consistent challenge in NLP development involving Yorùbá and under low-resource languages. Studies have also shown the presence of quality issues for existing corpora and models [93], which limits the reliability of findings from using such corpora. Consequently, efforts towards creating readily available corpora for specific NLP tasks are sacrosanct since better results mostly require high-quality and large data [44].
@@
Other language tools, including the syntactic parser and POS tagger, have also been reported to mostly fail in accounting for Yorùbá language linguistic nuances [45], thereby constituting a challenge requiring the development of language-specific tools for these specific tasks. Similarly, an essential NLP task such as effective spell-checking development in Yorùbá is still at the early development [37] due to the non-consideration of diacritic necessity in earlier studies [92], resulting in their limitation. A limited number of studies [54, 97, 37] involving computational NLP have been seen to include Yorùbá language in these domains. Furthermore, the majority of Yorùbá texts are also seen to be typed using plain ASCII, without diacritics [101], perhaps due to limited tools supporting their full implementation. This situation also induces lower quality translation between Yorùbá language and most European languages due to a lack of an adequate elision resolution tool [16]. Thus, there is a growing need for flexible multilingual pre-processing solutions.
Cultural and Societal Factors: An African language, such as Yorùbá, is not only a mirror into the mind of its users but also a mirror into their culture and history [91]. This emphasises the richness of the culture and language. However, younger generations of Yorùbá speakers increasingly favour English for formal and digital communication, while the parents also decide not to teach their infants many times [1]. This shift reduces the volume of contemporary Yorùbá texts and contributes to the decline of the language in digital contexts. Moreover, despite the evident progress in MT tasks for low-resource languages, NMT models still lag in accurately carrying out automatic translation involving cultural references [13] for Yorùbá language. Consequently, cultural and societal challenges of Yorùbá in NLP development require considerable attention, as they could potentially endanger the language in the face of westernization, globalization, and inter-ethnic communication [91].
@@ 5 Discussion
5.1 General Overview
This systematic review was intended to capture the growth and current stage of NLP participation for Yorùbá language over a decade. Through the information synthesised from the primary studies, it is obvious that significant progress has been made in notable areas of NLP involving Yorùbá language. MT has particularly received more attention through the inclusion of the language primarily in bilingual MT research [108, 47, 112, 61, 6, 46, 57] involving translation between two languages. Related studies investigating the inherent prerequisite towards enhancing translation qualities in MT model qualities, the development of MT tools, such as vowel elision resolution [16], and the application of rough set theory [62] in translation, have also been explored. Moreover, the need to create better quality data [93] has also led towards improved research in this domain through the development of additional parallel corpora and leveraging multilingual language pre-training to improve the model training capability in the presence of large high-quality datasets.
Similar to MT research, significant efforts have been made in TTS synthesis, with considerable efforts towards corpus development, as evidenced by publicly available datasets, most of which are speech corpora. Even though it is the primary focus in most research investigating the task, some of the studies have included ASR and speech-to-text [90, 29, 22] in their experiments, contributing to the entire research. Furthermore, sentiment analysis and information retrieval tasks involving Yorùbá have also benefited significantly from multilingual corpus development through the need to develop language resources for low-resource African languages. Most have been through the works in [85] and [19].
@@ 5.1 General Overview
Furthermore, recent studies have showcased the importance of multilingual and cross-lingual techniques, as they help to promote language availability for many African languages simultaneously. This is evident from named entity recognition [10, 12], sentiment analysis [86, 85, 111], and information retrieval tasks [89, 19, 128]. Also, models pre-trained on multilingual datasets have been seen to be beneficial for a low-resource language like Yorùbá. Overall, it is also evident that collaborative, open-source and community initiatives have greatly improved research and availability of resources in the bid to overcome resource scarcity whilst fostering knowledge-sharing among these resource-scarce African languages.
@@
However, in the context of constraints stunting the rapid growth of NLP in Yorùbá language, studies have highlighted mainly linguistic complexity such as diacritic dependency [101] and tonal variation [67]. Other challenges involving cultural and societal factors were also highlighted, including the primary hindrance, which has always been a limitation in the availability of corpora or the quality of such available corpora.
Ultimately, the geographical distribution of lead authors for the studies shows promising insight into the research efforts towards promoting Yorùbá in NLP involvement is global and not restricted to the primary home of the language. It also revealed necessary collaborations among authors and its pivot towards collectively improving resource availability for African languages.
This synthesis of findings not only elucidates the current state of Yorùbá NLP but also offers broader lessons for the development of NLP in other low-resource contexts. In many respects, the linguistic and infrastructural challenges identified here are mirrored in numerous other low-resource languages (LRLs). This makes it possible to extend the methodological insights gained from Yorùbá NLP to a wider set of languages with similar constraints.
5.2 Implications for Other Low-Resource Languages
Although this review focuses on the Yorùbá language, many of the observed trends, challenges, and methodological solutions are highly transferable to other LRLs, particularly those with similar linguistic complexity, socio-cultural context, or infrastructural limitations. Challenges such as dearth of annotated corpora, limited pre-trained language models, and tonal complexity are not unique to Yorùbá; they also hinder progress in languages like Igbo, Hausa, Wolof, Amharic, and Somali [101, 81].
From task NLP tasks perspective, diacritic restoration, which was addressed for Yorùbá language through neural network–based approaches [101, 44, 35] using datasets like the Lagos-NWU Yorùbá Speech Corpus could be adapted to other tone-marked or diacritic-dependent languages such as Ewe, Twi, or Vietnamese, as seen in the publications. Similarly, machine translation pipelines that combined rule-based, statistical, and neural methods [2, 59, 6, 67, 47, 61, 80, 5] and resources like MENYO-20k [13] and IkiniYorùbá [27] could be replicated for other LRLs with parallel corpora, enabling both bilingual and multilingual translation systems. Furthermore, in sentiment analysis tasks, multilingual resources such as AfriSenti [85] and NaijaSenti [86] have shown that transfer learning and multilingual pre-training can deliver competitive performance even for code-mixed text; this is an approach directly applicable to languages with high social media usage and code-switching patterns.
Moreover, from a methodological perspective, techniques such as transfer learning from high-resource to low-resource contexts [54, 89, 50, 93, 100, 95, 3, 63, 107], multilingual pre-training with mBERT, XLM-R, among others [10, 54, 50, 93], and cross-lingual embeddings as in MasakhaNER [10] represent broadly applicable strategies. These methods allow resource-constrained languages to leverage shared morphosyntactic or semantic structures in related languages. Furthermore, community-driven speech corpus development initiatives such as ÌròyìnSpeech [90] demonstrate how volunteer-based, open-licence data collection can bootstrap Automatic Speech Recognition (ASR) and Text-to-Speech (TTS) systems for other low-resource languages with oral traditions.
Furthermore, socio-cultural factors identified in Yorùbá NLP, such as code-switching [86], [16], domain-specific vocabulary gaps [80], [28], and reduced digital use [65], [40], are also common in many other low-resource languages. Addressing these requires culturally informed NLP design, specifically incorporating language-specific orthography. In addition, as shown in Yorùbá NLP, integrating such considerations into dataset curation and model fine-tuning not only improves performance but also enhances user adoption and long-term language dynamism. Although cultural and societal barriers, such as the preference for English in formal language usage, might superficially be interpreted as a decline in relevance, it is logically seen instead as evidence of a critical gap in technological inclusion, based on available research [1]. Yorùbá language remains a vibrant, widely spoken language, and its strong cultural presence highlights that the challenge is not diminishing demand, but rather insufficient technological support. Thus, the limitation of NLP tools for Yorùbá should be understood as a call for urgent resource development, ensuring the language remains visible and usable in digital technology, rather than as an indicator of declining societal importance.
In summary, the Yorùbá NLP landscape offers a transferable blueprint for advancing NLP in other low-resource languages through the evident foundational resources, application of hybrid methodologies, adopting multi-lingual pre-training and cross-lingual transfer, thereby accelerating progress, and the incorporation of socio-cultural knowledge into the technology lifecycle. These principles, grounded in the evidence presented in Tables LABEL:tab5 - LABEL:tab7, provide a replicable framework for fostering inclusive and sustainable NLP development across diverse low-resource language contexts. By considering the case of the Yorùbá language within a broader low-resource languages landscape, this discussion emphasizes the potential for synergy of methods, tools, and collaborative frameworks, thereby helping to ensure that research findings from one language can be effectively leveraged to support many others.
5.3 Current Yorùbá NLP in Relation to State-of-the-Art Practices
Existing studies in Yorùbá NLP reveal both progress and significant gaps when compared to mainstream NLP research. In terms of task coverage, work has primarily focused on foundational areas such as part-of-speech tagging, machine translation, sentiment analysis, and limited speech recognition. However, several high-impact tasks, including question answering, dialogue systems, summarization, sarcasm detection, identification of harmful language usage, including abusive, offensive, hate speech, or cyberbullying, and also multimodal applications, remain largely unexplored.
Methodologically, Yorùbá NLP has mostly relied on rule-based systems and statistical and traditional machine learning approaches [127, 95, 45], with limited research tasks leveraging transformer-based architectures and PLMs. In contrast, state-of-the-art NLP is dominated by large-scale pretrained models like Grok, GPT, LLaMA, Qwen models, among others, and task-specific fine-tuning, which enable broad generalization and high accuracy. The relative scarcity of Yorùbá-specific training data in these domains has limited the application of these methods.
Furthermore, while some corpora, lexicons, and parallel datasets exist as presented in Table LABEL:tab7, they are often fragmented or domain-limited. Unlike high-resource languages, Yorùbá NLP lacks many benchmark datasets and evaluation suites that facilitate reproducibility and comparison across approaches.
Altogether, these gaps highlight clear priorities for future research, by looking into the current reality of NLP involving Yorùbá, and the general state-of-the-art. Efforts should focus on developing benchmark datasets for core tasks, adopting more advanced transfer learning and multilingual fine-tuning strategies, and expanding into overlooked task areas such as dialogue systems and multimodal NLP.
5.4 Limitations of Study